Two types of electromechanical carburetors are being assembled and tested. Each of the first type requires 17 minutes of assembly time and 9 minutes of testing time. Each of the second type requires 13 minutes of assembly time and 8 minutes of testing time. If 324 minutes of assembly time and 175 minutes of testing time are available, how many of the second type can be assembled and tested if all the time is used?Express your answer correct to three significant digits and do not include units in your answer.
Answer by Theo(13342)

x = number of first type
y = number of second type

your constraint equations are:
17x + 13y = 324
9x + 8y = 175

the first equation is for assembly time.
the second equation is for testing time.

solve these two equations simultaneously to get:
x = 16.68421053
y = 3.105263158

if all the available time has to be used, then 3.105263158 of the second type can be assembled and tested.

the second type would use 13 * y = 40.36842105 hours of assembly time and 8 * y = 24.84210526 hours of testing time.

a graphical solution to solving those two equations simultaneously is shown below:

$$$

the graphical solution is the intersection of the two lines formed by the equations.